OpenStudy (anonymous):

What is the slope of the line through points p( 1,2) and q(7,-3)

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

I got -6/5

zepdrix (zepdrix):

Slope is given by, \[\Large \frac{\Delta y}{\Delta x} \qquad=\qquad \frac{y_2-y_1}{x_2-x_1}\] Looks like we were given,\[\Large (x_1,y_1)=(1,2) \qquad\qquad\qquad (x_2,y_2)=(7,-3)\] Plugging everything in gives us,\[\Large \frac{-3-2}{7-1}\] Hmmm I think maybe you have it upside down XD

OpenStudy (anonymous):

Oh okay! i got it. There is another similar question, What is the equation of the line with the slope of 3 and through the point 4 , 1

zepdrix (zepdrix):

The equation of a line in slope-intercept form is, \[\Large y=mx+b\]Where the slope is our \(\Large m\) and the y-intercept is \(\Large b\) So we can plug the slope value directly in for \(\Large m\). Then to solve for \(\Large b\), we'll have to plug in the coordinate pair they gave us,\[\Large y=3x+b\]Plug in your \(\large (4,1)\) to find b! :)

OpenStudy (anonymous):

What do you mean by plug in? where? @zepdrix

zepdrix (zepdrix):

\[\Large (x,y)=(4,1)\]Plug 4 in for x, 1 in for y. Into this,\[\Large y=3x+b\]

OpenStudy (anonymous):

y=12+1

OpenStudy (anonymous):

no no wait

OpenStudy (anonymous):

1 = 12 + b? @zepdrix

zepdrix (zepdrix):

mm ya that looks right, so what do you get for b? solve for b! :O

OpenStudy (anonymous):

-11

zepdrix (zepdrix):

Ok good. Now we want to back up a step, leave the (4,1) out of the equation and plug in your m and b to get the final form we're looking for.

OpenStudy (anonymous):

y = 3x -11 ?

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

okay one last problem I promise!!

OpenStudy (anonymous):

Solve the system of equations 3x + 3y = 6 5x - 6y =15

zepdrix (zepdrix):

There are two really convenient methods for solving linear systems like this, `substitution` and `elimination`. Do either of those sound familiar? :O

OpenStudy (anonymous):

Ive used subsitution in the past, but it wont seem to work right now

zepdrix (zepdrix):

It would work, it would just take a few extra steps. Elimination might be a little easier for this one. :)

zepdrix (zepdrix):

We want to match up the coefficients on `one` of our variables in both equations. In this equation let's try to match up the coefficient on our x's. After we do that, we can subtract the equations and the x's will cancel out. That will allow us to solve for y. Hmm let's multiply this first equation by 5,\[\Large 3x+3y=6 \qquad\to\qquad 15x+15y=30\]

zepdrix (zepdrix):

And then we want to multiply our second equation by 3, both sides as we did before,\[\Large 5x-6y=15 \qquad\to\qquad ?\]How bout you do this one ^^

OpenStudy (anonymous):

Wait, how do we know what to multiply by

OpenStudy (anonymous):

i got it though, 15x -18y = 45

OpenStudy (anonymous):

and I understand how we know what to multiply by

zepdrix (zepdrix):

Oh you figured that out? :) k cool.

OpenStudy (anonymous):

yea, let me see if I can do the rest

zepdrix (zepdrix):

\[\Large 15x+15y=30\]\[\Large 15x -18y = 45\]k :)

OpenStudy (anonymous):

x = 35/70

OpenStudy (anonymous):

i mean 75/30

OpenStudy (anonymous):

x = 2.5

zepdrix (zepdrix):

you solved for x first? That's strange :o we wanted to solve for y first. That's why we matched up the values on x. So they would cancel out when we subtract the equations.

OpenStudy (anonymous):

y = -25

zepdrix (zepdrix):

Hmmm not quite :O Let's see... subtracting the bottom equation from the top one gives us,\[\Large 0x+33y=-15\]

OpenStudy (anonymous):

how 33y?

OpenStudy (anonymous):

15 - 18

zepdrix (zepdrix):

We're subtracting this equation from the other one, \[\Large 15x -18y = 45\] So we have to remember that the subtraction applies to EVERY term in this equation,\[\Large -(15x -18y) = -(45)\]

zepdrix (zepdrix):

So it becomes a positive 18 when we distribute that negative sign.

OpenStudy (anonymous):

why would we distribute a - sign

zepdrix (zepdrix):

That's what we do when we subtract multiple terms. We have to apply the subtraction to each term. So it flips the sign of everything.

OpenStudy (anonymous):

were subtracting? I thought we were adding

zepdrix (zepdrix):

\[\Large 15x+15y=30\]\[\Large 15x -18y = 45\] If we added these equations together, it would give us 30x :[ that's not what we want. We want to cancel out the x's.

OpenStudy (anonymous):

so we multiply each by -

zepdrix (zepdrix):

yes :) maybe that's a better way to explain it hehe

OpenStudy (anonymous):

so it changes them to 0x -15y = -30 0x +18 = -45

zepdrix (zepdrix):

ah woops, i shoulda been more clear :( we want to multiply `one of the equations` by -1.

OpenStudy (anonymous):

so were doing the bottom one? we get 0x + 15y = 30 0x + 18 = -45

zepdrix (zepdrix):

The way you wrote those 0's a little strange :) We want to perform all of the addition/subtraction at the same time. But yes it looks good.